OpenStudy (anonymous):

|-19| + 32 = 80y - |-31| - 79y

OpenStudy (anonymous):

y=82

OpenStudy (anonymous):

thats wrong

OpenStudy (anonymous):

19+32= 80y -31-79y (take absolute value of 19 and 31. 31 will still be negative because of subtraction sign outside of absolute value) 51 = y-31 (add 19 and 32. subtract 80y and 79y) 82=y (add 31 to both sides) and thats it:)